# The effect of self-induced Marangoni flow on polar-nematic waves in active-matter systems

### v1

@@ 5 Conclusion
The dynamics of a dry system with local nematic alignment is described by a truncated Boltzmann equation derived in Ref. [14, 13]. Here we have adapted the model for wet systems by incorporatinhg an advection term that results from the Marangoni flow. After establishing a simplified dynamical model that couples density, polar order parameter and nematic order parameter fields, we have focused of the case of a long-wave approximation where the local flow field is proportional to the local density gradient. To characterise the degree of departure from the dry regime, we have introduced a Marangoni parameter that measures the relative strength of the Marangoni flow as compared to the self-propulsion speed of the agents.

### v2

@@ 5 Conclusion
The dynamics of a dry system with local nematic alignment is described by a truncated Boltzmann equation derived in Ref. [14, 13]. Here we have adapted the model for wet systems by incorporating an advection term that results from the Marangoni flow. After establishing a simplified dynamical model that couples density, polar order parameter and nematic order parameter fields, we have focused of the case of a long-wave approximation where the local flow field is proportional to the local density gradient. To characterise the degree of departure from the dry regime, we have introduced a Marangoni parameter that measures the relative strength of the Marangoni flow as compared to the self-propulsion speed of the agents.
@@
Author contribution statement
All authors contributed equally to conceptualisation, writing and literature research. (AP) numerical simulations, bifurcation analysis and figure preparation. (UT) figure preparation and analysis.
@@ Acknowledgements
[…]
Data availability All data generated or analysed during this study are included in this published article [and its supplementary information files]. Source codes are available from AP on request.
Appendix A Closed system of hydrodynamic equations for the leading order parameter fields in the presence of advective flow
Consider the two-dimensional motion of self-propelled agents in the presence of an advection flow field $\mathbf{U}(c)$ which only depends on the local concentration of the agents $c$ and is insensitive to their local orientation. Following [12, 13, 14] we write the Boltzmann equation (1) in the angular Fourier space in terms of the order parameter fields $\rho_{q}$ defined by equation (4)
$\partial_{t}\rho_{q}+\mathbf{\nabla}\cdot(\mathbf{U}\rho_{q})+\frac{v_{0}}{2}(\partial\rho_{q-1}+\partial^{*}\rho_{q+1})=d\partial\partial^{*}\rho+R_{q}\rho_{q}+\sum_{n=-\infty}^{\infty}I_{n,q}\rho_{n}\rho_{q-n}.$ (24)
Here, $\partial=\partial_{x}+i\partial_{y}$, $R_{q}<0$ describes random re-orientation due to rotational diffusion and $I_{n,q}$ is the collision matrix in the angular Fourier space. Note that $R_{q}$ and $I_{n,q}$ depend on microscopic parameters such as the interaction distance, the alignment rule during two-particle collisions and the statistics of rotational diffusion. For example, as shown in [14], $R_{q}=-\lambda(1-e^{-(q\sigma^{2}/2)})$, where $\lambda$ is the rate at which the agents randomly change their orientation from $\theta$ to $\theta+\eta$, where $\eta$ is normally distributed with zero mean and standard deviation $\sigma$. The collisions matrix $I_{n,q}$ consists of two parts: the first part describes polar alignment between two agents that collide at an acute angle and the second part corresponds to antipolar alignment when the collision occurs at an obtuse angle. For simplicity of discussion, we do not show the expression for $I_{n,q}$ in terms of the specific rules for collision statistics.
Assuming $\rho_{q>4}=0$, we obtain for $\rho_{3,4}$ the kinetic equations
$\begin{split}\partial_{t}\rho_{3}+\mathbf{\nabla}\cdot(\mathbf{U}\rho_{3})+\frac{v_{0}}{2}(\partial\rho_{2}+\partial^{*}\rho_{4})&=d\mathbf{\nabla}^{2}\rho_{3}+R_{3}\rho_{3}+\sum_{n=-1}^{4}I_{n,3}\rho_{n}\rho_{3-n},\\ \partial_{t}\rho_{4}+\mathbf{\nabla}\cdot(\mathbf{U}\rho_{4})+\frac{v_{0}}{2}(\partial\rho_{3})&=d\mathbf{\nabla}^{2}\rho_{4}+R_{4}\rho_{4}+\sum_{n=0}^{4}I_{n,4}\rho_{n}\rho_{4-n}.\end{split}$ (25)
After applying the scaling (5) to equations (25) we retain only terms up to the order $\epsilon^{3}$. At this order, the concentration $c$ is replaced with its mean value $c_{0}$. The relationship between $\rho_{4}$ and $(c,\rho_{1,2})$ is strictly local, while $\rho_{3}$ also depends on the gradient of $\rho_{2}$, namely,
$\rho_{4}=-\frac{I_{2,4}\rho_{2}^{2}}{R_{4}+c_{0}(I_{0,4}+I_{4,4})},~~\rho_{3}=\frac{(v_{0}/2)\partial\rho_{2}-I_{1,3}\rho_{1}\rho_{2}}{R_{3}+c_{0}(I_{0,3}+I_{3,3})}.$ (26)
Note that the closure relations (26) do not depend on the advecting surface flow $\mathbf{U}$, and therefore exactly coincide with those derived in Refs. [13, 14]. In consequence, the resulting closed system of hydrodynamic equations for $c$, $\rho_{1,2}$ for a wet system correspond to the hydrodynamic equations derived in Refs. [13, 14] each extended by the advection term $\mathbf{\nabla}\cdot(\mathbf{U}\rho_{q})$ with $(q=0,1,2)$ corresponding to the self-induced Marangoni flow.